# Filter Like You Test: Data-Driven Data Filtering for CLIP Pretraining

### v2

@@ 1 Introduction
We introduce methods that together achieve state-of-the-art performance on the DataComp medium scale filtering benchmark [17]. At the core is Filter Like You Test [^1] The name is inspired by FLYP [18]. (FLYT), a method for training a scoring model that predicts the weight an image-text pair should have in CLIP pretraining. Mixing-FLYT (M-FLYT) is an implementation of FLYT that uses scores generated by existing scoring methods as input features to the scoring model. Complementing FLYT, Soft Cap Sampling (SCS) is a simple sampling strategy that leverages the probabilistic structure of FLYT’s scores to construct the final pretraining dataset.
@@
1 Input: Scoring model ${q}_{{\phi}_{0}}$ with initial parameters ${\phi}_{0}$, reference model ${f}_{{\theta}_{0}}$ with initial parameters ${\theta}_{0}$, feature extractor ${\Psi}$, upstream dataset ${D}$, downstream dataset ${C}$
2 Parameters: Number of steps $T$, upstream data batch size ${B}$, downstream data batch size ${B^{\prime}}$, weighted loss ${\ell_{\mathrm{WCL}}}$, and downstream loss ${\ell_{\mathrm{DS}}}$
3 for $t=0$ to $T-1$ do
@@ 1 Input: Training examples ${z}_{1:{M}}$ and their scores ${s}_{1:{M}}$
2 Parameters: Repetition penalty ${\alpha}$, target dataset size ${N}$ and sample count per iteration ${G}$
3 ${\hat{D}}=\left[\text{ }\right]$
@@ 5       ${w}_{1:{M}}=\mathrm{softmax}({s}_{1:{M}})$
6       $i_{1:{G}}=$ Sample ${G}$ indices according to distribution ${w}_{1:{M}}$ without repetitions.
@@ 2 Related work
Ren et al. [37] introduced an algorithm for online example weighting which utilizes gradient signals from a validation dataset for producing per-example weights. Shu et al. [43] employs a similar idea to Ren et al. [37] with the addition of a weighting model that is optimized to weigh the examples. FLYT is similar to these methods in that it too learns weights and backpropogates through a “target loss” after a gradient update on the “training loss.” However, unlike FLYT, these methods consider online data weighting for standard supervised learning, with the intention of improving robustness to noise and data imbalance, mostly in small-scale settings.
@@ 3 Method
In this section, we describe FLYT in detail. We introduce our notation by going over standard CLIP training (Section 3.1), then move to describing FLYT (Section 3.2). We then go over different feature extractors for FLYT’s implementation (Section 3.3), followed by our choice of downstream data (Section 3.4), and explaining the weighted CLIP loss we use to train FLYT (Section 3.5). Finally, we present SCS (Section 3.6).
@@ 3.1 Notation and problem setting
We denote a CLIP training example ${z}=({{z}^{I}},{{z}^{T}})\in{D}$ where ${D}$ is our pretraining dataset and ${{z}^{I}},{{z}^{T}}$ are the training image and text, and denote a batch of ${B}$ training examples by ${{z}_{1:{B}}}$. We train a CLIP model with parameters ${\theta}$ that maps each example ${z}$ to features ${{f}_{{\theta}}({z})}=({{f}_{{\theta}}({{z}^{I}})},{{f}_{{\theta}}({{z}^{T}})})$. Letting ${{f}_{{\theta}}({{z}_{1:{B}}})}$ denote the features of all the examples in the batch, the standard CLIP loss is
@@ 3.2 Filter Like You Test (FLYT)
In this section we describe FLYT. Our goal is to learn a scoring model ${q}_{\phi}$ parameterized by ${\phi}$. The scoring model takes a feature vector ${\zeta}={\Psi}({z})$ produced by applying feature extractor ${\Psi}$ to example ${z}$, and maps it to a score ${s}$, such that the higher ${s}$ is, the more useful ${z}$ is as a pretraining example. We overload notation to let
@@
Next, we introduce the key part of our approach, a batch of downstream data (e.g., the ImageNet training set). Let ${{\hat{z}}_{1:{B^{\prime}}}}$ denote this batch (note that ${B^{\prime}}$ does not necessarily have to be equal to ${B}$). We use this batch to obtain a gradient signal for updating ${{w}_{1:{B}}}$ and hence the scoring model parameters ${\phi}$. To that end, define:
@@
3.3 Feature extractors
Mixing-FLYT (M-FLYT)
Inspired by past success in mixing multiple filtering heuristics [17, 22, 49, 53], we implement FLYT using existing scoring methods as our feature extractor. Specifically, the feature extractor ${\Psi}$ produces a vector ${\zeta}\in\mathbb{R}^{{k}}$ of ${k}$ different scores, generated by ${k}$ different scoring methods. Our scoring model ${q}_{\phi}:\mathbb{R}^{{k}}\to\mathbb{R}$ then takes as input the different scores to output a single unified score ${s}$. We refer to this implementation of a scoring model as a mixing model.
Embedding-FLYT (E-FLYT)
With E-FLYT, we use a pretrained CLIP model as the feature extractor ${\Psi}{}$ to generate image and text embeddings. Our scoring model ${q}_{\phi}$ then maps the concatenated image and text embeddings to a score ${s}$.
3.4 Choice of downstream data
@@ 3.6 Soft Cap Sampling (SCS)
A straightforward approach would be to sample from this distribution 128M times to create our final dataset. In practice, doing that results in over-represented examples. To correct this, we introduce SCS (Soft Cap Sampling): every time we add an example to our filtered data, we subtract a “repetition penalty” $\alpha$ from its score, reducing the probability it is sampled again. For implementation efficiency, we perform the sampling procedure iteratively. In each iteration, we sample a batch of ${G}=100K$ examples without repetitions, then subtract ${\alpha}$ from the score of each batch member. Algorithm 2 presents SCS, and Figure 2 illustrates the distribution generated via SCS.
@@ 4 Experiments
Table 1: Comparison of our method to top performing methods on the DataComp medium scale filtering benchmark. DFN-FT is our best reproduction of DFN [16]. * denotes results from the original paper, and $+$ indicates evaluations using our downloaded 119M candidate pool using the authors’ provided UIDs. Table 14 compares our evaluations to the results reported in the corresponding papers.
@@ 4.2 M-FLYT results
HYPE ${-d_{L}}$ 0.45
HYPE ${\epsilon_{i}}$0.05
@@
IN-weighted ${r}=2$ 0.350 0.364
IN-weighted ${r}=4$ 0.354 0.359
IN-weighted ${r}=8$ 0.353 0.365
IN-weighted ${r}=16$ 0.352 0.362
@@
[…]
We compare M-FLYT to its individual input components and to several baseline approaches. These baselines include a simple sum of all input scores and a sum of scores after standardization to zero mean and unit standard deviation. Finding that standardization significantly improves performance, we further explore weighted standardized sums where inputs are weighted by their normalized ImageNet scores adjusted to achieve a specific max-to-min weight ratio ${r}$, for details see Section C.5.
@@ 4.3 SCS results
For SCS, we experiment with a repetition penalty ${\alpha}$ between 0.1 and 0.6, and sample in batches of ${G}=100K$. We choose ${G}$ to balance efficient sampling while ensuring the resulting dataset closely approximates what we would get using ${G}=1$. Section C.10 shows that the exact value of ${G}$ has minimal impact on performance. When preparing the filtered dataset, we use a 1.5M shuffle buffer in the DataComp resharder to prevent batches from containing many copies of the same example.
@@ 4.4 Small scale results
We evaluate our method on the small scale DataComp benchmark. We use the mixing model trained by M-FLYT described in Section 4.2, and sample using SCS with a repetition penalty ${\alpha}=0.5$ and a batch size ${G}=10K$. Table 3 shows that our method achieves 8% zero-shot ImageNet zero-shot accuracy and 18.5% average accuracy, outperforming previous work on both categories. These results suggest that M-FLYT and SCS can be effectively used across different computational scales. Figure 4 shows the effect of different ${\alpha}$ values.
@@ 5 Limitations and future work
Another notable limitation is that E-FLYT underperforms, though its scores did contribute to M-FLYT. To surpass state-of-the-art methods we had to leverage existing filtering methods through M-FLYT rather than relying on a general feature extractor. We believe this is likely a deficiency in our optimization stack, solvable by more careful tuning of learning rate (schedules) for individual components and regularization techniques such as dropout. In similar vein, we believe that more careful parameter tuning should allow us to leverage downstream data that is more diverse than ImageNet, though these hypotheses require empirical validation.
A third limitation—and direction for future research—is that we use data weighting as a proxy for data selection when defining the FLYT objective function. It would be interesting to try to model data selection more directly via reinforcement learning: at each reference model training step, the scoring model picks a batch of training data and observes a reward signal for the chosen data.
A fourth direction for future work is closing the data selection / pretraining loop by iteratively re-training the scoring model to select the next data to pretrain our model on. FLYT is particularly well-suited to this form of active learning, as it allows us to use the pretraining model as the reference model at each scoring iteration, ensuring each selected batch is optimal for the model’s current state.
@@ Acknowledgments
We thank Sonia Lichtenzveig Sela and Alexander Tsverin from the TAU CS systems team for crucial technical support. We thank Alex Fang and Vaishaal Shankar for helpful discussion of the DFN implementation details, and Sanghyuk Chun for sharing the HYPE codebase.
This research was partially supported by the computing resources of the Tel Aviv University Center for AI and Data Science, and the Israeli Science Foundation (ISF) grant no. 2486/21.
@@ Appendix A Implementing data-parallel distribution
where ${f}_{i}$ is shorthand for the embedding ${f}_{{\theta}}({z}_{i})$ that implicitly says we only need to store the value of the embedding and not the computational graph behind it. This way each GPU calculates ${g}$ for its own sub-batch, and we can then gather and sum their result to get ${g}$ for the full batch using the CLIP gradient accumulation technique from Cherti, Mehdi et al. [7]. Using this, we define
@@
which we can again compute using gradient accumulation. Treating ${v}$ as a constant vector independent of ${{w}_{1:{B}}}$, the gradient we would like to calculate is
@@ Appendix B M-FLYT input details
    HYPE. Kim et al. [22] provided a PyTorch implementation which includes their hyperbolic CLIP weights, and the functionality required to reproduce the image specificity ${\epsilon_{i}}$, text specificity ${\epsilon_{t}}$, and negative Lorentzian distance ${-d_{L}}$. We used these three values as inputs.
@@
[…]
@@ B.1 DFN reproduction
Since M-FLYT requires good inputs to perform well, we made effort to train a high performing DFN. In the paper, Fang et al. [16] trained a ViT-B/32 CLIP model initialized with OpenAI’s checkpoint on the high-quality private dataset HQITP-350M. Since we do not have access to this dataset, we trained our DFN on CC12M [41]. Similar to Fang et al. [16], we then fine-tune on the combined training sets of IN1K with sampled OpenAI templates as texts, MS COCO, and Flickr30k. Unless otherwise specified, we use the same hyper-parameters as DataComp medium scale training. Tables 4 and 5 show results we obtained with different hyperparameter choices when our reproductions DFN-Base and DFN-FT, respectively.

### v3

@@ 1 Introduction
We consider the setting of the DataComp filtering benchmark [17]. Our goal is to find a filtering method that selects a subset from a candidate pool of web-sourced image-text training data. To evaluate the quality of the selected subset, it is fed to a fixed training pipeline and the resulting model is evaluated on a fixed set of zero-shot image comprehension tasks. The DataComp rules do not constrain the computational cost of the filtering method. Moreover, DataComp allows using external data (except evaluation test sets) for making the filtering decisions, so long as the selected training set is a subset of the candidate pool.
In this setting, we introduce methods that together achieve state-of-the-art performance on the DataComp medium scale. At the core is Filter Like You Test [^1] The name is inspired by FLYP [18]. (FLYT), a method for training a scoring model that predicts the weight an image-text pair should have in CLIP pretraining. Mixing-FLYT (M-FLYT) is an implementation of FLYT that uses scores generated by existing scoring methods as input features to the scoring model. Complementing FLYT, Soft Cap Sampling (SCS) is a simple sampling strategy that leverages the probabilistic structure of FLYT’s scores to construct the final pretraining dataset.
@@
1 Input: Scoring model ${q}_{{\phi}_{0}}$ with initial parameters ${\phi}_{0}$, reference model ${f}_{{\theta}_{0}}$ with initial parameters ${\theta}_{0}$, feature extractor $\Psi$}, upstream dataset D{D}, downstream dataset C{C}
2 Parameters: Number of steps $T$, upstream data batch size B{B}, downstream data batch size $B^{\prime}$}, weighted loss $\ell_{\mathrm{WCL}}$}, and downstream loss $\ell_{\mathrm{DS}}$}
3 for $t=0$ to T1T-1 do
@@ 1 Input: Training examples ${z}_{1:{M}}$ and their scores ${s}_{1:{M}}$
2 Parameters: Repetition penalty $\alpha$}, target dataset size N{N} and sample count per iteration G{G}
[…]
@@ 5 ${w}_{1:{M}}=\mathrm{softmax}({s}_{1:{M}})$
6 $i_{1:{G}}=$ Sample G{G} indices according to distribution ${w}_{1:{M}}$ without repetitions.
@@ 2 Related work
FLYT shares conceptual and technical similarities with meta-learning paradigms. Particularly similar meta-learning works address the problem of learned data weighting. Ren et al. [37] introduced an algorithm for online example weighting which utilizes gradient signals from a validation dataset for producing per-example weights. Shu et al. [43] employs a similar idea to Ren et al. [37] with the addition of a weighting model that is optimized to weigh the examples. FLYT is similar to these methods in that it too learns weights and backpropogates through a “target loss” after a gradient update on the “training loss.” However, unlike FLYT, these methods consider online data weighting for standard supervised learning, with the intention of improving robustness to noise and data imbalance, mostly in small-scale settings.
@@ 3 Method
In this section, we describe FLYT in detail. We introduce our notation by going over standard CLIP training (Section 3.1), then move to describing FLYT (Section 3.2). We then go over the weighted CLIP loss we use to train FLYT (Section 3.3), followed by two different feature extractors for FLYT’s implementation (Section 3.4), and our choice of downstream data (Section 3.5). Finally, we present SCS (Section 3.6).
@@ 3.1 Notation and problem setting
We denote a CLIP training example ${z}=({{z}^{I}},{{z}^{T}})\in{D}$ where D{D} is our pretraining dataset and ${{z}^{I}},{{z}^{T}}$ are the training image and text, and denote a batch of B{B} training examples by ${z}_{1:{B}}$}. We train a CLIP model with parameters $\theta$} that maps each example z{z} to features ${{f}_{{\theta}}({z})}=({{f}_{{\theta}}({{z}^{I}})},{{f}_{{\theta}}({{z}^{T}})})$. Letting ${f}_{{\theta}}({{z}_{1:{B}}})$} denote the features of all the examples in the batch, the standard CLIP loss is
@@ 3.2 Filter Like You Test (FLYT)
In this section we describe FLYT. Our goal is to learn a scoring model ${q}_{\phi}$ parameterized by $\phi$}. The scoring model takes a feature vector ${\zeta}={\Psi}({z})$ produced by applying feature extractor $\Psi$} to example z{z}, and maps it to a score s{s}, such that the higher s{s} is, the more useful z{z} is as a pretraining example. We overload notation to let
@@
Next, we introduce the key part of our approach, a batch of downstream data (e.g., the ImageNet training set). Let ${\hat{z}}_{1:{B^{\prime}}}$} denote this batch (note that $B^{\prime}$} does not necessarily have to be equal to B{B}). We use this batch to obtain a gradient signal for updating ${w}_{1:{B}}$} and hence the scoring model parameters $\phi$}. To that end, define:
@@ 3.3 Weighted CLIP loss
3.4 Feature extractors
Mixing-FLYT (M-FLYT)
Inspired by past success in mixing multiple filtering heuristics [17, 22, 49, 53], we implement FLYT using existing scoring methods as our feature extractor. Specifically, the feature extractor $\Psi$} produces a vector ${\zeta}\in\mathbb{R}^{{k}}$ of k{k} different scores, generated by k{k} different scoring methods. Our scoring model ${q}_{\phi}:\mathbb{R}^{{k}}\to\mathbb{R}$ then takes as input the different scores to output a single unified score s{s}. We refer to this implementation of a scoring model as a mixing model.
Embedding-FLYT (E-FLYT)
With E-FLYT, we use a pretrained CLIP model as the feature extractor $\Psi$}{} to generate image and text embeddings. Our scoring model ${q}_{\phi}$ then maps the concatenated image and text embeddings to a score s{s}.
3.5 Choice of downstream data
@@ 3.6 Soft Cap Sampling (SCS)
A straightforward approach would be to sample from this distribution 128M times to create our final dataset. In practice, doing that results in over-represented examples. To correct this, we introduce SCS (Soft Cap Sampling): every time we add an example to our filtered data, we subtract a “repetition penalty” $\alpha$ from its score, reducing the probability it is sampled again. For implementation efficiency, we perform the sampling procedure iteratively. In each iteration, we sample a batch of G=100K{G}=100K examples without repetitions, then subtract $\alpha$} from the score of each batch member. Algorithm 2 presents SCS, and Figure 2 illustrates the distribution generated via SCS.
@@ 4 Experiments
Table 1: Comparison of our method to top performing methods on the DataComp medium scale filtering benchmark. DFN-FT is our best reproduction of DFN [16]. * denotes results from the original paper, and ++ indicates evaluations using our downloaded 119M candidate pool using the authors’ provided UIDs. Table 15 compares our evaluations to the results reported in the corresponding papers.
@@ 4.2 M-FLYT results
HYPE $d_{L}$} 0.45
HYPE $\epsilon_{i}$} -0.05
@@
IN-weighted r=2{r}=2 0.350 0.364
IN-weighted r=4{r}=4 0.354 0.359
IN-weighted r=8{r}=8 0.353 0.365
IN-weighted r=16{r}=16 0.352 0.362
@@
We create and reproduce 12 input scores. We use the CLIP score of OpenAI’s ViT-B/32 and ViT-L/14 models [35], as well as two reproduced DFNs [16], before and after fine-tuning, which we refer to as DFN-Base and DFN-FT (Section B.1). We reproduce the image and text specificity ${\epsilon_{i}}{},{\epsilon_{t}}{}$, and negative Lorentzian distance $d_{L}$}{} introduced by HYPE [22] using their codebase, and the s-CLIPLoss and $\mathrm{{NormSim}}_{\infty}(\mathrm{{IN1K}})$}{} scores introduced by Wang et al. [49] using theirs. We also add three new inputs. Two binary classifiers that classify between some “high quality” dataset and the DataComp medium scale dataset, and an E-FLYT scoring model described in Section C.7. See Appendix B for more details about the inputs.
We compare M-FLYT to its individual input components and to several baseline approaches. These baselines include a simple sum of all input scores and a sum of scores after standardization to zero mean and unit standard deviation. Finding that standardization significantly improves performance, we further explore weighted standardized sums where inputs are weighted by their normalized ImageNet scores adjusted to achieve a specific max-to-min weight ratio r{r}, for details see Section C.5.
@@
In Appendix E we complement our quantitative results with a qualitative comparison of filtering methods. In particular, we plot the top-scoring training example according to different filtering methods. Figures 5, 6 and 7 show that M-FLYT prioritizes examples containing multiple ImageNet classes, while sorting by OpenAI’s ViT-B/32 CLIP scores favors images with visible text that matches the caption text.
@@ 4.3 SCS results
For SCS, we experiment with a repetition penalty $\alpha$} between 0.1 and 0.6, and sample in batches of G=100K{G}=100K. We choose G{G} to balance efficient sampling while ensuring the resulting dataset closely approximates what we would get using G=1{G}=1. Section C.10 shows that the exact value of G{G} has minimal impact on performance. When preparing the filtered dataset, we use a 1.5M shuffle buffer in the DataComp resharder to prevent batches from containing many copies of the same example.
@@
Table 3: Using SCS with M-FLYT compared to IN-weighted.
Mixing method Sampling method ImageNet Average
M-FLYT Threshold top 20% 0.359 0.371
M-FLYT SCS (${\alpha}=0.15$) 0.401 0.377
IN-weighted Threshold top 20% 0.353 0.365
IN-weighted SCS (${\alpha}=0.15$) 0.379 0.362
To test SCS independently of M-FLYT, we apply it with a repetition penalty ${\alpha}=0.15$ to scores from our strongest baseline, IN-weighted with $r=8$. We convert these scores to probabilities via softmax, following the same approach as with M-FLYT. Table 3 shows that SCS improves IN-weighted ImageNet accuracy by 2.6% while decreasing average score by 0.3%. In comparison, applying SCS to M-FLYT improves ImageNet accuracy by 4.2% and average score by 0.6%. This demonstrates that a key advantage of FLYT is that it naturally produces a distribution over training examples, which can be leveraged through sampling approaches like SCS.
@@ 4.4 Small scale results
We evaluate our method on the small scale DataComp benchmark. We use the mixing model trained by M-FLYT described in Section 4.2, and sample using SCS with a repetition penalty ${\alpha}=0.5$ and a batch size G=10K{G}=10K. Table 4 shows that our method achieves 8% zero-shot ImageNet zero-shot accuracy and 18.5% average accuracy, outperforming previous work on both categories. These results suggest that M-FLYT and SCS can be effectively used across different computational scales. Figure 4 shows the effect of different $\alpha$} values.
@@ 5 Limitations and future work
The computational overhead of FLYT is a potential limitation, but may also be negligible in some scaling regimes. In our experiments, M-FLYT achieves high performance using approximately $3\times$ fewer FLOPs than training a DataComp medium-scale model, due to training on relatively few examples (20M). Scaling up the number of training steps would proportionally increase computational costs, as detailed in Section C.2. Whether the scoring model needs to increase in computational cost as the data and final model size increases—and hence the relative overhead of FLYT in large scale settings—remains a topic for future work.
Another notable limitation is that E-FLYT underperforms methods such as DFN, though its scores did contribute to M-FLYT. To surpass state-of-the-art methods we had to leverage existing filtering methods through M-FLYT rather than relying on a general feature extractor. We believe this is likely a deficiency in our optimization stack, solvable by more careful tuning of learning rate (schedules) for individual components and regularization techniques such as dropout. In similar vein, we believe that more careful parameter tuning should allow us to leverage downstream data that is more diverse than ImageNet, though these hypotheses require empirical validation.
A fourth limitation—and direction for future research—is that we use data weighting as a proxy for data selection when defining the FLYT objective function. It would be interesting to try to model data selection more directly via reinforcement learning: at each reference model training step, the scoring model picks a batch of training data and observes a reward signal for the chosen data.
A fifth direction for future work is closing the data selection / pretraining loop by iteratively re-training the scoring model to select the next data to pretrain our model on. FLYT is particularly well-suited to this form of active learning, as it allows us to use the pretraining model as the reference model at each scoring iteration, ensuring each selected batch is optimal for the model’s current state.
@@ Acknowledgments
We thank Sonia Lichtenzveig Sela and Alexander Tsverin from the TAU CS systems team for crucial technical support. We thank our anonymous NeurIPS 2025 reviewers for experiment and writing suggestions, Alex Fang and Vaishaal Shankar for helpful discussion of the DFN implementation details, and Sanghyuk Chun for sharing the HYPE codebase.
This research was partially supported by the computing resources of the Tel Aviv University Center for AI and Data Science, the Adelis Foundation, and the Israeli Science Foundation (ISF) grant no. 2486/21.
@@ Appendix A Implementing data-parallel distribution
where ${f}_{i}$ is shorthand for the embedding ${f}_{{\theta}}({z}_{i})$ that implicitly says we only need to store the value of the embedding and not the computational graph behind it. This way each GPU calculates g{g} for its own sub-batch, and we can then gather and sum their result to get g{g} for the full batch using the CLIP gradient accumulation technique from Cherti, Mehdi et al. [7]. Using this, we define
@@
which we can again compute using gradient accumulation. Treating v{v} as a constant vector independent of ${w}_{1:{B}}$}, the gradient we would like to calculate is
@@ Appendix B M-FLYT input details
    HYPE. Kim et al. [22] provided a PyTorch implementation which includes their hyperbolic CLIP weights, and the functionality required to reproduce the image specificity $\epsilon_{i}$}, text specificity $\epsilon_{t}$}, and negative Lorentzian distance $d_{L}$}. We used these three values as inputs.
@@
    s-CLIPLoss. Wang et al. [49] also provided a PyTorch implementation. We used OpenAI ViT-B/32 embeddings to calculate s-CLIPLoss and $\mathrm{{NormSim}}_{\infty}(\mathrm{{IN1K}})$}.
@@ B.1 DFN reproduction
Since M-FLYT requires good inputs to perform well, we made effort to train a high performing DFN. In the paper, Fang et al. [16] trained a ViT-B/32 CLIP model initialized with OpenAI’s checkpoint on the high-quality private dataset HQITP-350M. Since we do not have access to this dataset, we trained our DFN on CC12M [41]. Similar to Fang et al. [16], we then fine-tune on the combined training sets of IN1K with sampled OpenAI templates as texts, MS COCO, and Flickr30k. Unless otherwise specified, we use the same hyper-parameters as DataComp medium scale training. Tables 5 and 6 show results we obtained with different hyperparameter choices when our reproductions DFN-Base and DFN-FT, respectively.